Proton Polarization at Room Temperature

in collaboration with the laboratory for Laser Spectroscopy in Kyoto Univ.

The polarized target are the important tools in particle and nuclear physics. Usually, the dynamic nuclear polarization method(DNP) is used for the polarization of protons in organic materials containing a few percents of complex[1]. The electron polarization is transfered to protons by means of irradiation of microwave in a high magnetic field(> 2.5T) at a temperature lower than 1 K.

We are trying to polarize protons in an organic crystal by laser excitation, and subsequent dynamic polarization of protons on an excied triplet state. The idea has been proposed by Henstra, et al. at Leiden, recently[2]. Our sample is the single crystal of Naphthalene doped with Pentacene(0.01mol%). Electron pairs in a Pentacene molecule are diamagnetic in the ground state, because these electrons are on a singlet state(S=0). Electrons are excited to a higher singlet state by N_{2}-laser(λ = 3.37nm). The spin-orbital interaction causes the transition from the singlet excited state to the intermidiate states which are triplet(S=1). If the magnetic field(H ∼ 3.3kG) is applied along the X-axis of Pentacene(Figure), electrons on the intermidiate state are aligned. The populations of the triplet state S_{x}=+1,0,−1 is 12%, 76%, and 12%, respectively. The life time of the triplet state is about 20μsec. Then the polarization is transfered from electrons to protons in Pentacene and Naphthalene by means of microwave irradiation (f ∼ 10GHz) during the life time of this intermidiate state.

After the transition of electrons to the ground state, the proton spin is kept polarized for long time, since the spin–spin interaction between electrons and protons is negligible.

Figure 1: Energy Levels of electrons (pentacene)

We got the proton polarization 1000 times as high as the one in the thermal equilibrium condition. It corresponds to about 0.1% polarization. Recently we have carried out the experiment in lower temperature than 0C^{∘} in order to obtain higher polarization. It has been found that the relaxation time of the proton spin in the crystal depends upon the temperature in the temperature region higher than –20C^{∘}. It is about 40min. at room temperature. It becomes much shorter due to slight heating of the crystal by the laser power. At liquid nitrogen temperature, the relaxation time was found to be about 1000min. The polarization is about 13% at this temperature. We expect higher polarization by using other types of laser. Furthermore we are studying the possiblity of higer polarization in higer temperature than liquid nitrogen temperature. We believe that it will be applicable to the polarized proton target.

References

[1] A.Masaike, H.Glattli, J.Ezratty and A.Malinovski, Phys. Lett. 30A (1969) 63

[2] A.Henstra, T-S.Lin, J.Schumidt and W.Th.Wenckbach, Chem. Phys. Letters 165 (1990) 6